### v1

# Zero-Shot EEG-to-Gait Decoding via Phase-Aware Representation Learning

Abstract
Accurate decoding of lower-limb motion from EEG signals is essential for advancing brain–computer interface (BCI) applications in movement intent recognition and control. However, challenges persist in achieving causal, phase-consistent predictions and in modeling both inter- and intra-subject variability. To address these issues, we propose NeuroDyGait, a domain-generalizable EEG-to-motion decoding framework that leverages structured contrastive representation learning and relational domain modeling. The proposed method employs relative contrastive learning to achieve semantic alignment between EEG and motion embeddings. Furthermore, a multi-cycle gait reconstruction objective is introduced to enforce temporal coherence and maintain biomechanical consistency. To promote inter-session generalization, during fine-tuning, a domain dynamic decoding mechanism adaptively assigns session-specific prediction heads and learns to mix their outputs based on inter-session relationships. NeuroDyGait enables zero-shot motion prediction for unseen individuals without requiring adaptation and achieves superior performance in cross-subject gait decoding on benchmark datasets. Additionally, it demonstrates strong phase-detection capabilities even without explicit phase supervision during training. These findings highlight the potential of relational domain learning in enabling scalable, target-free deployment of BCIs.
{IEEEkeywords}
domain generalization, electroencephalography, contrastive learning, gait decoding, zero-shot prediction
1 Introduction
Restoring human mobility using brain-computer interfaces (BCIs) is an emerging focus in neurorehabilitation [colucci2022brain]. A key component of this effort is the decoding of lower-limb motor intent from non-invasive electroencephalography (EEG) signals [he2018brain]. This capability is essential for enabling assistive technologies such as exoskeletons and neuroprosthetics [kilicarslan2021neuro]. Although recent progress has demonstrated the feasibility of EEG-based gait decoding, significant challenges remain for real-world deployment. A major limitation lies in the limited availability of normal motion data from patients with neurological impairments. These individuals are often unable to perform standard gait patterns. This makes it critical to design decoders that generalize across individuals without relying on patient-specific fine-tuning, thereby enabling practical and scalable use in rehabilitation settings [azab2019weighted, fahimi2019inter].
Despite this progress, two key challenges hinder reliable decoding of gait dynamics. The first limitation concerns the widespread adoption of segment-to-point prediction strategies. These methods use a short window of historical EEG signals to estimate the motion state at a single final time step [he2018mobile, brantley2018full, shafiul2020prediction, presacco2011neural].This formulation satisfies the requirements of causal decoding. However, it reduces human motion to isolated time points, neglecting the temporal continuity and biomechanical constraints that characterize natural gait. In practice, the context of gait provides critical information for inferring joint angles, as joint configurations are strongly influenced by surrounding temporal dynamics. Human locomotion is intrinsically structured and phase-dependent: even under similar motor intent, joint trajectories can vary significantly across individuals, and identical joint configurations may reflect different gait phases depending on context and prior states [hamacher2017exploring, presacco2012decoding]. Without explicitly modeling this structure, predictions risk becoming unstable or semantically inconsistent. Therefore, a robust decoding framework must incorporate these underlying dynamics to enable the model to differentiate between gait states that are superficially similar but functionally distinct across different cycles.
The second challenge lies in the limitations of current cross-domain learning strategies, which often treat multiple sessions or subjects as independent sources. These methods typically use shared encoders for generalization or subject-specific heads for individual differences [jayaram2016transfer, chen2021ms], but fail to capture structured relationships across domains. This oversimplification becomes particularly problematic when generalizing across heterogeneous data sources, where domain shifts—caused by factors such as sensor placement, individual differences, and recording conditions—introduce structured dependencies between domains. Empirical evidence indicates that both inter-session and inter-subject variability can substantially degrade the consistency and reliability of biosignal decoding. This highlights the importance of models that not only capture domain-specific characteristics but also exploit shared structure across domains to enable more robust and transferable representation learning [huang2023discrepancy, maswanganyi2022statistical, ng2023deep]. Therefore, a robust framework should be designed to model both intra-domain specificity and inter-domain structure to enable more adaptive and generalizable learning.
To address these challenges, we propose NeuroDyGait, a domain-generalizable EEG-to-gait decoding framework built on semantically structured representation learning and relational domain modeling. The name reflects the model’s focus on neural (Neuro) dynamics (Dy) underlying continuous human locomotion (Gait), emphasizing its ability to capture temporally evolving brain-motor relationships for robust cross-subject prediction. Our contributions are twofold:
  • Phase-Aware Pretraining with Relative Contrastive Learning: We propose a novel pretraining strategy based on relative contrastive learning that aligns EEG and motor embeddings by modeling sample-wise semantic similarity. Unlike traditional contrastive approaches that rely on predefined positive-negative pairs, our method employs a learnable distance function to infer relative similarity among all samples within a batch. This relative similarity modeling approach enables the model to learn fine-grained, phase-aware gait semantics and distinguish between motion states that are functionally distinct but visually similar. Furthermore, the model is trained to reconstruct multi-cycle gait trajectories from EEG embeddings, which helps it capture neuromusculoskeletal dynamics in context and enhances its robustness to noisy or atypical inputs.
  • Relation-Aware Decoding via Domain Head Mixture: We propose a domain-aware decoding mechanism that assigns a dedicated prediction head to each source session. During fine-tuning, a learnable domain head dynamically integrates intra-domain outputs with a weighted combination of cross-domain predictions, enabling the model to capture structured inter-session dependencies. This adaptive fusion of domain-specific knowledge enhances generalization across subjects by leveraging latent relationships among heterogeneous signal patterns, thereby enabling more robust and context-aware decoding in real-world scenarios.
2 Related Work
2.1 EEG-Based Motor Execution Decoding
EEG-based neural decoding of lower-limb motor execution has been widely studied, leveraging EEG’s high temporal resolution for capturing motor control dynamics. Pfurtscheller and colleagues [pfurtscheller2001motor] demonstrated that event-related desynchronization (ERD) and synchronization (ERS) effectively characterize motor execution and imagery, laying the groundwork for decoding lower-limb movements. Ang and colleagues [ang2012filter] used common spatial pattern (CSP) and linear discriminant analysis (LDA) to decode gait phases, enhancing motor-related EEG feature extraction. Schirrmeister et al. [schirrmeister2017deep] employed convolutional neural networks (CNNs), boosting accuracy in classifying lower-limb motor imagery. Recent deep learning methods have further improved spatiotemporal feature extraction from EEG signals [goh2018spatio, tortora2020deep, fu2022matn], demonstrating their ability to capture complex nonstationary neural patterns. Finally, Wang et al. [wang2018implementation] developed a real-time BCI system for lower-limb exoskeleton control, validating the feasibility of EEG-driven motor decoding in rehabilitation settings. While these works significantly advance EEG-based BCIs, they typically focus on classifying discrete movement states or rely on pointwise regression, often lacking biomechanical awareness and dynamic consistency in continuous motion decoding.
2.2 Cross-Subject Generalization in EEG Decoding
A major challenge in EEG decoding lies in inter-subject variability, where differences in brain dynamics, head geometry, and sensor placement cause distributional shifts between individuals. To address this, domain adaptation methods have been applied to align source and target feature distributions through adversarial training [li2018cross], statistical moment matching [long2015learning], or subspace projection [liang2022multi]. However, most approaches assume access to target-domain data during training, which is often unrealistic in clinical settings where normative data from impaired users are unavailable. Domain generalization methods aim to overcome this by learning domain-invariant representations solely from source data [wang2024dmmr, lu2023hybrid], yet often treat all source domains independently or equally. In practice, EEG and motor signals vary both across and within subjects, and ignoring inter-session relationships can limit generalization.
In this work, we model session-level structure explicitly by assigning each source session a dedicated decoder and learning a dynamic mixture of their outputs. This relation-aware strategy captures both intra-session specificity and cross-session similarity, improving robustness in zero-shot subject transfer.
2.3 Contrastive Representation Learning for Structured Movement
Contrastive learning has proven effective for self-supervised representation learning in structured movement tasks like gait analysis and motor imagery decoding. By comparing similarities and differences between samples, it captures discriminative features without needing labels.
In video-based action recognition, motion-aware frameworks such as MaCLR align visual and motion modalities for enhanced video representations [xiao2022maclr], while MCL emphasizes motion through alignment of gradient maps and optical flow [li2021motion]. For skeleton-based action recognition, contrastive learning captures structural dynamics [guo2022contrastive], with HiCLR enforcing consistency across hierarchical augmentations [zhang2023hierarchical], and cross-modality approaches modeling complex patterns [li2023cross].
In EEG-based motor imagery, contrastive learning addresses inter-subject variability using CNNs and attention mechanisms [li2024self], while supervised contrastive learning improves gait recognition using EEG and EMG signals [fu2023gait].
These developments highlight the strength of contrastive learning in modeling dynamic, structured movement across various modalities.
3 Methods
3.1 Overview
Our proposed framework, NeuroDyGait, consists of two training stages: (1) a pretraining stage (see Stage I in Fig.1) that learns temporally structured and semantically meaningful EEG embeddings by reconstructing synchronized motion signals and modeling cross-domain relationships, optimized using a combination of reconstruction, contrastive, and prediction losses; and (2) a domain generalization stage (see Stage II in Fig.1) that employs a session-wise head architecture with a learnable domain fusion mechanism, optimized via a domain fusion loss and a supervised prediction loss.
Figure 1: Overview of NeuroDyGait architecture. Stage I: The dual-encoder model consists of an EEG encoder, a motor encoder, and a decoder. It is trained with reconstruction, contrastive, and prediction losses to extract biomechanically meaningful EEG representations. Stage II: During training, session-specific heads and a domain weighting layer are optimized to predict final motor state. Test: The model computes a normalized mixture of all source-domain heads for zero-shot generalization.
An overview of NeuroDyGait, including the Stage I, Stage II, and Test Stage, is shown in Fig. 1.
3.2 Stage I: Pretraining with Biomechanical Supervision
3.2.1 Dual-Encoder Architecture
The pretraining stage employs two modality-specific encoders: an EEG encoder $f_{e}$ and a motor encoder $f_{m}$. EEG signals within 2-second windows are processed by $f_{e}$ to produce embeddings $\mathbf{z}_{e}=f_{e}(\mathbf{x})$, while synchronized joint angle sequences are passed through $f_{m}$ to produce $\mathbf{z}_{m}=f_{m}(\mathbf{y})$. A motor decoder $g$ reconstructs the motion signal $\hat{\mathbf{y}}=g(\mathbf{z}_{e})$ from the EEG embeddings.
The EEG encoder is composed of a deep convolutional feature extractor that captures multi-scale spatiotemporal patterns across channels, followed by a final constrained $1\times T$ convolution layer to project the output into a compact latent space. The use of filter growth (25, 50, 100, 200) across layers promotes hierarchical abstraction, and norm-constrained weights ensure stability in training.
The motor encoder combines temporal convolutional blocks with a Transformer module. Specifically, stacked 1D convolutions reduce temporal resolution while enriching local features, and a Transformer encoder equipped with positional encoding captures long-range temporal dependencies in the joint trajectories. The resulting sequence representation is aggregated via average pooling to obtain a fixed-length latent embedding.
The motor decoder $g$ reconstructs time-domain joint angle sequences $\hat{\mathbf{y}}=g(\mathbf{z}_{e})$ from EEG embeddings using a transposed CNN architecture. It first projects the latent vector into a spatiotemporal tensor, then applies a series of transposed convolutional layers with batch normalization and ReLU activations. The final layer outputs 6-dimensional joint angles, and dynamic output padding ensures the sequence length matches the original (400 time steps). This structure supports end-to-end training and preserves temporal continuity in motion reconstruction.
3.2.2 Reconstruction Loss
To ensure the EEG embedding captures biomechanically meaningful information, we train the decoder to reconstruct the full 2-second motion sequence from EEG embeddings using mean squared error (MSE):
$\mathcal{L}_{\text{rec}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}\|^{2}$
where $N$ denotes the number of training samples in a batch.
This loss encourages the EEG representation to retain sufficient information for accurately capturing motor dynamics over time.
3.2.3 Prediction Loss
While the reconstruction loss encourages the EEG embeddings to preserve the full temporal dynamics of motion, our ultimate objective is to accurately predict the gait state at the final time point of the sequence. To this end, we introduce a prediction loss that focuses explicitly on the final frame of the motor output:
$\mathcal{L}_{\text{pred}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{i,T}-\mathbf{y}_{i,T}\|^{2}$
where $\hat{\mathbf{y}}_{i,T}$ and $\mathbf{y}_{i,T}$ denote the predicted and ground truth motor states at the final time step $T$ for the $i$-th sample, and $N$ is the number of training samples in a batch.
This loss encourages the EEG embedding to preserve information that is most relevant for predicting the motor state at the last frame, thereby reducing the potential influence of earlier, less relevant gait phases during pretraining.
3.2.4 Relative Contrastive Learning
Inspired by Xu et al.  [xu2024relcon], we introduce a relative contrastive learning framework that captures fine-grained semantic relationships between EEG and motor signals. Unlike traditional contrastive methods that define fixed positive and negative sets, our approach treats all intra-batch samples as candidates and enforces relative similarity rankings between them. This encourages the model to learn an embedding space where the distance structure reflects the semantic alignment between modalities.
Cross-Attention-Based Distance
Given an EEG embedding $\mathbf{z}_{e}\in\mathbb{R}^{d}$ and a candidate motor embedding $\mathbf{z}_{m}\in\mathbb{R}^{d}$, we define a cross-attention mechanism that treats $\mathbf{z}_{e}$ as the query and $\mathbf{z}_{m}$ as the key and value. The attention weight is computed via a scaled dot-product:
$\alpha=\text{softmax}\left(\frac{\langle W_{q}\mathbf{z}_{e},W_{k}\mathbf{z}_{m}\rangle}{\sqrt{d}}\right)$ (1)
The attended motor feature is computed as:
$\hat{\mathbf{z}}_{\text{m}}=W_{o}(\alpha\cdot W_{v}\mathbf{z}_{m})$ (2)
where $W_{o}\in\mathbb{R}^{d\times d}$ are learnable projection matrices. $W_{o}$ maps the aggregated attention output back to the shared latent space. This allows the attended motor representation $\hat{\mathbf{z}}_{\text{m}}$ to be directly compared to the EEG embedding $\mathbf{z}_{\text{e}}$ for distance computation, enabling cross-modal alignment in a unified embedding space.
The cross-modal distance is measured as the mean squared error between the EEG embedding and the attended motor representation:
$d(\mathbf{z}_{e},\mathbf{z}_{m})=\left|\hat{\mathbf{z}}_{m}-\mathbf{z}_{e}\right|_{2}^{2}$ (3)
This learned distance function serves as a semantic similarity measure between EEG and motor embeddings, capturing alignment based on shared motion-related features.
Relative Contrastive Loss
Using the pairwise distances $d(\mathbf{z}_{e}^{(i)},\mathbf{z}_{m}^{(j)})$ between all EEG-motor pairs in a batch, we define a relative contrastive loss that enforces ranking consistency. For each anchor $i$, we compute the similarity matrix:
$S_{ij}=-\frac{d(\mathbf{z}_{e}^{(i)},\mathbf{z}_{m}^{(j)})}{\tau}+b$ (4)
where $\tau$ is a temperature scaling parameter, and $b$ is a learnable scalar bias. For each anchor, we sort all candidates by similarity and compute a listwise loss based on log-softmax over the sorted similarities:
$\mathcal{L}_{i}=-\sum_{j=1}^{B-1}\left(S_{ij}-\log\sum_{k=j}^{B-1}e^{S_{ik}}\right)$ (5)
Here, we exclude the self-similarity term $S_{ii}$ to avoid trivial comparisons, resulting in a total of $B-1$ candidates for each anchor in the loss computation.
The final relative contrastive loss is averaged across all anchors:
$\mathcal{L}_{\text{rcl}}=\frac{1}{B}\sum_{i=1}^{B}\mathcal{L}_{i}$ (6)
This loss encourages the model to assign higher similarity to more semantically aligned motor sequences while preserving the relative ranking among all candidates.
The total loss of stage I is a weighted sum of all components:
$\mathcal{L}_{\text{stage I}}=\mathcal{L}_{\text{rec}}+\lambda_{\text{pred}}\mathcal{L}_{\text{pred}}+\lambda_{\text{rcl}}\mathcal{L}_{\text{rcl}}$
We set $\lambda_{\text{pred}}=0.5$ to moderately encourage accurate motor prediction during pretraining without dominating the learning dynamics, allowing the model to prioritize representation alignment through the relative contrastive loss. Meanwhile, $\lambda_{\text{rcl}}=1$ emphasizes the preservation of semantic consistency in EEG-motor relations, which is crucial for robust downstream decoding.
3.3 Stage II: Prediction with Domain Relation-Aware Decoding
3.3.1 Session-Specific Prediction Heads
In the second stage, we retain only the pretrained EEG encoder and attach a prediction head to each source-domain session. For a given EEG embedding $\mathbf{z}_{e}$ belonging to session $s$, the corresponding session-specific head $h_{s}$ outputs a prediction $\hat{\mathbf{y}}=h_{s}(\mathbf{z}_{e})$ for the final motor frame (i.e., the gait pattern at the end of the 2-second window). Each session head is trained using the mean squared error (MSE) loss:
$\mathcal{L}_{\text{sup}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}\|^{2}$
3.3.2 Domain Weighting Layer and Mixture Prediction
To leverage inter-session relationships during inference, we introduce a domain weighting mechanism that adaptively fuses predictions from all source-domain heads except the one corresponding to the input session [yaoimproving]. Given an EEG embedding $\mathbf{z}_{e}\in\mathbb{R}^{d}$ from session $s$, the mixture prediction is computed as:
$\hat{\mathbf{y}}_{\text{mix}}=\sum_{k=1}^{N_{\text{src}}}\alpha_{k}(\mathbf{z}_{e})\cdot h_{k}(\mathbf{z}_{e})$
where $h_{k}(\cdot)$ denotes the domain-specific head for session $k$, and $\alpha_{k}(\mathbf{z}_{e})\in[0,1]$ is the attention weight predicted by a lightweight scoring network:
$\boldsymbol{\alpha}=\text{softmax}\left(W_{a}\mathbf{z}_{e}+\mathbf{b}_{a}+\log(1-\mathbf{m})\right)$
Here, $W_{a}\in\mathbb{R}^{N_{\text{src}}\times d}$ and $\mathbf{b}_{a}\in\mathbb{R}^{N_{\text{src}}}$ are learnable parameters, and $\mathbf{m}\in\{0,1\}^{N_{\text{src}}}$ is a one-hot mask where $m_{s}=1$ denotes the current session. The additive masking term $\log(1-\mathbf{m})$ sets the attention logit of the input session to $\infty$- ∞, ensuring that $\alpha_{s}=0$. This excludes the intra-domain head from the mixture, promoting the learning of a cross-domain feature aggregation function.
3.3.3 Joint Optimization with $\mathcal{L}_{\text{sup}}$ and $\mathcal{L}_{\text{df}}$
We jointly train the session-specific heads and the domain weighting mechanism using two objectives: - $\mathcal{L}_{\text{sup}}$: ensures each session head captures reliable subject-specific gait mapping. - $\mathcal{L}_{\text{df}}$: ensures the mixture prediction $\hat{\mathbf{y}}_{\text{mix}}$ is close to ground truth:
$\mathcal{L}_{\text{df}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{\text{mix},i}-\mathbf{y}_{i}\|^{2}$
The total loss of stage II is a weighted sum of supervised Loss and domain fusion loss:
$\mathcal{L}_{\text{stage II}}=\mathcal{L}_{\text{sup}}+\lambda_{\text{df}}\mathcal{L}_{\text{df}}$
We set $\lambda_{\text{df}}=1$ to equally balance the supervised objective and the domain fusion objective, encouraging the model to maintain high prediction accuracy within each session while simultaneously learning to generalize across domains through adaptive head mixing.
3.3.4 Inference via Head Mixture
During inference on an unseen domain, we discard all session identity information and compute the final prediction as a weighted combination over all available source-domain heads:
$\hat{\mathbf{y}}_{\text{test}}=\sum_{i=1}^{N_{\text{src}}}\alpha_{i}\cdot h_{i}(\mathbf{z}_{e}^{tgt})$
where weights $\alpha_{i}$ are predicted by the domain weighting layer. This strategy enables zero-shot generalization by leveraging learned inter-session relationships without requiring any target-domain adaptation.
4 Experiment
4.1 Dataset
4.1.1 Gait-EEG Dataset
To investigate the brain mechanisms involved in walking, we collected a new dataset, Gait-EEG Dataset (GED) [fu2025eeg2gait], recording brain activity along with simultaneous lower-limb joint angles natural walk on level ground. The dataset contains the recordings from 50 able-bodied participants (25 males, 25 females; aged 21 to 46, mean age 28.4, standard deviation 5.2), with no history of neurological disorders or lower limb pathologies. Participants engaged in two independent level-ground walking experiment sessions, with every session comprising three identical walking blocks. Each block included approximately 40 trials, with each trial representing EEG signals and synchronized lower-limb joint angles as the participant walked straight for 7.7 meters. Sessions were spaced at least three days apart. The dataset includes synchronized recordings from a 60-channel active EEG, a 4-channel electrooculogram (EOG), along with measurements from six joint angle sensors (bilateral hips, knees, and ankles) [fu2025eeg2gait].
This study has been reviewed and approved by the Institutional Review Board (IRB-2021-709) of Nanyang Technological University, ensuring compliance with applicable legislation, ethical and safety requirements in Singapore. All participants have provided informed consents before the experiment.
4.1.2 Open-access Dataset
To further validate our proposed method, we conducted additional experiments using the open-access Full Body Mobile (FBM) brain-body imaging dataset [brantley2018full]. The FBM dataset contains full-body motion capture data (66 markers) from approximately 10 walking trials performed by 10 able-bodied individuals under various gait conditions, including level ground, ramp, and stair walking. Data were collected using a wireless IMU-based motion capture system, enabling natural, unconstrained movement. To characterize lower-limb motor states, we selected the Z-axis joint angles of eight bilateral joints: hip, knee, ankle, and ball of the foot. A data segmentation stride of 50 ms was used.
4.2 Data Preprocessing
EEG signals from both datasets were preprocessed using an identical pipeline. A bandpass filter ranging from 0.1 to 48 Hz was first applied to eliminate low-frequency drifts and high-frequency noise. The filtered signals were then re-referenced using common average referencing (CAR) to reduce spatially correlated noise across channels. Subsequently, the EEG signals were resampled to 200 Hz to reduce computational overhead while preserving relevant neural activity. Similarly, goniometer signals were resampled to 200 Hz and normalized to zero mean and unit variance on a per-joint basis. Channels with zero standard deviation were safely handled by substituting a value of one to avoid division errors.
4.3 Data Segregation
4.3.1 Gait-EEG Dataset
For NeuroDyGait, which is designed for cross-subject generalization, we split the 50 participants in the Gait-EEG dataset into training, validation, and test sets in an 8:1:1 ratio. Specifically, data from participants 1–40 were used for training, participants 41–45 for validation, and participants 46–50 for testing. All available data from each participant were included in their respective splits.
4.3.2 FBM Dataset
Similarly, in the FBM dataset, we used the complete data from participants 2–8 for training, participant 9 for validation, and participant 10 for testing. Data from participant 1 were excluded due to misalignment between the EEG and motion signals, which prevented reliable synchronization.
This subject-wise division ensures that the model is evaluated on entirely unseen individuals, emphasizing its generalization capability across subjects.
4.4 Evaluation Metric
We evaluated the efficacy of NeuroDyGait by comparing the predicted angles of joints with their actual recorded angles, using three standard regression metrics: Pearson correlation coefficient ($r$), coefficient of determination ($R^{2}$), and root mean squared error (RMSE). These metrics jointly capture trend consistency, explained variance, and absolute prediction accuracy.
$r=\frac{\operatorname{cov}(y,\hat{y})}{\sigma(y)\cdot\sigma(\hat{y})}$ (7)
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (8)
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (9)
Here, $y$ denotes the actual joint angle and $\hat{y}$ represents the predicted angle. The covariance between two variables $A$ and $B$ is denoted as $\operatorname{cov}(A,B)$, and $\sigma(A)$ is the standard deviation of $A$. $\bar{y}$ indicates the mean of the actual values. Each sequence represents data collected from a single trial over $n$ time steps, corresponding to $\frac{n}{100}$ seconds at a sampling rate of 100 Hz.
The Pearson $r$ value reflects the consistency in trend between the predicted and actual trajectories. The $R^{2}$ score measures the proportion of variance explained by the model, while RMSE captures the average magnitude of prediction error, penalizing larger deviations more heavily.
4.5 Implementation and Hyperparameter Settings
NeuroDyGait was implemented using the PyTorch library. Training was conducted using the Adam optimizer with default hyperparameter settings. A batch size of 512 was used, and training continued for a maximum of 50 epochs ($epoch_{\text{max}}$). Instead of using a fixed learning rate, we adopted a cosine learning rate scheduler with linear warm-up. The learning rate starts from zero and increases linearly during the initial warm-up phase, reaching the predefined maximum learning rate at the end of the warm-up period. After that, it gradually decreases to a minimum value following a cosine decay curve over the remaining training epochs. During the initial warm-up phase, spanning the first 2 epochs, the learning rate increases linearly from zero to the initial value. The total schedule is computed based on the number of training epochs and the number of iterations per epoch. The initial learning rate and minimum learning rate are listed in Table 1.
Table 1: Training hyperparameters used in the cosine learning rate schedule for each stage and dataset.
Setting Highest LR Min LR
GED (Stage I) 1e-3 1e-4
GED (Stage II) 2e-5 2e-6
FBM (Stage I) 1e-4 1e-5
FBM (Stage II) 5e-5 5e-6
5 Results and Analysis
In this section, we conduct a comprehensive evaluation of the proposed framework across multiple key dimensions. We begin by assessing model performance, comparing our method against several state-of-the-art baselines on two benchmark datasets: GED [fu2025eeg2gait] and FBM [brantley2018full]. The results demonstrate consistent improvements in motion prediction accuracy, highlighting the effectiveness of our approach. We then examine cross-dataset transferability, evaluating the model’s generalization capability across different datasets, which underscores its robustness and adaptability in out-of-distribution settings. To further probe the model’s internal behavior, we visualize phase-specific EEG embeddings using t-SNE, revealing structured, phase-aligned representations that reflect the temporal specificity of the learned features. Additionally, we analyze the relationship between domain attention entropy and out-of-domain predictive error, illustrating how the domain-aware decoding mechanism leverages inter-session dynamics to enhance decoding performance. We also perform spatial analysis via saliency mapping to identify cortical regions critical to the model’s predictions. The results indicate that the model predominantly focuses on central sensorimotor channels, such as Cz and CP2, consistent with the neural correlates of lower-limb motor control. Finally, we present a series of ablation studies to systematically assess the contribution of key architectural and training components to the overall performance of the proposed framework.
5.1 Model Performance
In this section, we validate the performance of the proposed method on GED and FBM dataset and compared with several state-of-the-art deep learning and machine learning algorithms in the BCI domain, including ContraWR [yang2021self], FFCL [li2022motor], TSception [ding2022tsception], Temporal Convolutional Network (TCN) [ingolfsson2020eeg], ST-Transformer [song2021transformer], EEGConformer [song2023eeg], SPaRCNet [jing2023development], EEGNet [lawhern2018eegnet], and deepConvNet [schirrmeister2017deep] and EEG2GAIT [fu2025eeg2gait].
The results of all metrics on the test sets are presented to evaluate the performance of NeuroDyGait in comparison to baseline methods on GED and FBM dataset. NeuroDyGait achieved an $r$ value of 0.673 ($R^{2}$ = 0.446, RMSE = 0.780) on GED dataset and an $r$ value of 0.259 ($R^{2}$ = 0.012, RMSE = 0.995) on FBM dataset, outperforming the baseline methods. All baseline methods were trained according to the strategy described in Section 4.3. A summary of the results for NeuroDyGait and the baseline methods on the different datasets is provided in Table 2 and 3. Across both GED and FBM datasets, NeuroDyGait consistently achieves either the best or second-best results on all evaluation metrics, demonstrating robust and competitive performance against existing state-of-the-art methods.
Table 2: The results of different methods on GED.
Method Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
ContraWR  [yang2021self] 0.2484 0.0515 1.0220
FFCL  [li2022motor] 0.3105 0.0743 1.0096
TSception  [ding2022tsception] 0.2783 0.0411 1.0271
EEGNet  [lawhern2018eegnet] 0.4450 0.1948 0.9413
TCN  [ingolfsson2020eeg] 0.2248 0.0448 1.0256
EEGConformer  [song2023eeg] 0.1279 -0.0099 1.0060
SPaRCNet  [jing2023development] 0.5946 0.3094 0.8713
ST-Transformer  [song2021transformer] 0.5764 0.3148 0.8679
deepConvNet  [schirrmeister2017deep] 0.6331 0.3928 0.8166
EEG2GAIT  [fu2025eeg2gait] 0.6348 0.3603 0.8385
NeuroDyGait 0.6725 0.4458 0.7799
↓: the lower the better; ↑: the higher the better.
The best results are highlighted in bold.
Table 3: The results of different methods on FBM.
Method Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
ContraWR  [yang2021self] 0.1315 -0.0926 1.0461
FFCL  [li2022motor] 0.1863 -0.0504 1.0259
TSception  [ding2022tsception] 0.1997 -0.0819 1.0409
EEGNet  [lawhern2018eegnet] $0$0.0622 1.0317
TCN  [ingolfsson2020eeg] 0.1348 -0.0304 1.0161
EEGConformer  [song2023eeg] 0.1279 -0.0099 1.0060
SPaRCNet  [jing2023development] 0.1347 -0.0669 1.0338
ST-Transformer  [song2021transformer] 0.1704 -0.0160 1.0111
deepConvNet  [schirrmeister2017deep] 0.2167 0.0063 0.9978
EEG2GAIT  [fu2025eeg2gait] 0.2189 0.0261 0.9879
NeuroDyGait 0.2585 0.0115 0.9952
↓: the lower the better; ↑: the higher the better.
The best results are highlighted in bold and the next best are marked using underlines.
5.2 Cross-Dataset Transferability
To evaluate the transferability of learned EEG representations, we conducted a cross-dataset experiment in which Stage I of NeuroDyGait was first pretrained on GED. The resulting EEG encoder was then paired with a randomly initialized motor encoder and decoder adapted to the FBM joint structure. This model served as the initialization for training on the FBM dataset, and was further trained through Stage I and Stage II using FBM data only.
This approach tests whether the EEG encoder, trained on one dataset, can provide a better initialization than random weights for a new domain, thereby facilitating the model in learning motor-relevant features and improving performance. Table 4 presents the results of this transfer setup compared to training NeuroDyGait from scratch on FBM. The transferred model achieves marginal improvements over training from scratch across all three metrics, suggesting better generalization and robustness of the EEG representations learned through cross-dataset pretraining.
Table 4: Cross-dataset transfer results (GED $\rightarrow$ FBM).
Init. Strategy Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
Scratch (FBM only) 0.2585 0.0115 0.9952
[…]
Initializing the EEG encoder from a Stage I model pretrained on GED improves performance on FBM, demonstrating cross-dataset transferability.
5.3 Enhancing Performance via Target Domain Fine-Tuning
Although NeuroDyGait outperforms or matches all baselines on both GED and FBM datasets, its predictive performance on FBM remains relatively modest (Pearson $r=0.259$, $R^{2}=0.012$), likely due to the small number of subjects and increased inter-subject variability.
To simulate realistic BCI deployment with limited calibration data, we performed target domain fine-tuning after training stage I and II by using only the first 3 minutes of EEG data per test session (less than 8% of total session length). The first 2.5 minutes were used for fine-tuning, and the remaining 30 seconds for early stopping validation.
As summarized in Table 5, fine-tuning substantially improved performance on both datasets, with particularly pronounced gains on FBM. On FBM, Pearson $r$ increased from 0.26 to 0.62—an improvement of over 2.4 times—and $R^{2}$ rose from 0.01 to 0.36, representing a more than 35-fold increase. Meanwhile, RMSE decreased by 0.19. On GED, fine-tuning also enhanced all metrics, increasing $r$ by 0.10 and reducing RMSE by 0.13.
Table 5: Effect of Fine-Tuning Using Only the First 3 Minutes of Test Session Data (<8% of Total Session Length)
Dataset Metric Original Fine-Tuned Improvement
GED Pearson $r$ ($\uparrow$) 0.6725 0.7748 +0.1023
$R^{2}$ Score ($\uparrow$) 0.4458 0.6005 +0.1547
RMSE ($\downarrow$) 0.7799 0.6480 -0.1319
FBM Pearson $r$ ($\uparrow$) 0.2585 0.6169 +0.3584
$R^{2}$ Score ($\uparrow$) 0.0115 0.3604 +0.3489
RMSE ($\downarrow$) 0.9952 0.8012 -0.1940
5.4 t-SNE Visualization of Phase-Specific EEG Embeddings
5.4.1 Kinematic Basis for Four‐Phase Segmentation
To inject gait‐cycle structure into our EEG embeddings, we partitioned each continuous gait cycle into four phases by detecting four reproducible kinematic events in the sagittal‐plane hip and knee trajectories of each leg [neptune2001contributions, anderson2004contributions, zajac2002biomechanics, whittle2014gait]:
  1. (a)
    Left hip maximal flexion (just prior to left toe‐off): marks end of left support and onset of left swing.
  2. (b)
    Left knee maximal flexion (mid‐swing peak): corresponds to peak elevation of the left limb for foot clearance.
  3. (c)
    Right hip maximal flexion (just prior to right toe‐off): marks end of right support and onset of right swing.
  4. (d)
    Right knee maximal flexion (mid‐swing peak): corresponds to peak elevation of the right limb.
By chaining these events in temporal order (a) → (b) → (c) → (d), we define four gait cycle phases:
  • Phase 1: From (a) to (b)
  • Phase 2: From (b) to (c)
  • Phase 3: From (c) to (d)
  • Phase 4: From (d) to the next (a)
Each phase reflects a distinct interlimb coordination pattern and captures transitions between swing and stance for both legs.
5.4.2 Visualization Method
Using the EEG encoder pretrained in Stage I, we extracted embeddings for overlapping EEG windows. Each window was labeled according to the phase of its final time‐sample kinematic phase and assigned a distinct color. We then projected all embeddings into two dimensions with t‐SNE [van2008visualizing], and overlaid the phase‐color labels using Matplotlib’s scatter() (alpha=0.7). As shown in Fig. 2, embeddings from different gait phases form spatially distinct clusters, demonstrating that the Phase‐Aware Pretraining objective implicitly captures the periodic structure of gait.
Figure 2: t-SNE visualization of EEG embeddings from before (upper panel) and after (lower panel) Stage I training. Colors represent different gait phases as defined in Section5.4.1.
5.5 Domain Attention Entropy and Predictive Performance
To better understand the behavior of our domain fusion mechanism, we analyzed the entropy of the attention weights produced by the domain attention module for each session in the test set. Specifically, we computed the entropy of the predicted domain weight distribution at inference time, where a higher entropy indicates broader reliance on multiple source domains, while a lower entropy reflects more selective attention to a few sources.
As shown in Fig. 3, we analyzed the relationship between attention entropy and decoding performance, measured by L1 Error. Each data point represents the mean entropy and L1 Error computed over a batch containing data from a single session. The results reveal a significant positive correlation between entropy and L1 Error, with a Pearson correlation coefficient (PCC) of 0.6673 and a $p\ll 1\times 10^{-5}$. The fitted curve, depicted in red, clearly illustrates this trend.
Figure 3: Scatter plot showing the relationship between domain attention entropy and L1 prediction error across test sessions.
5.6 Spatial Analysis
In addition to the performance metrics presented, we expanded our evaluation to include spatial feature importance analysis to better understand critical areas during decoding. To achieve this, we employed saliency mapping—a technique in machine learning that visualizes the importance of each input feature for the model’s predictions [simonyan2013deep]. This method highlights the input areas the model is most sensitive to when making predictions. The saliency map, $S$, is generated by calculating the gradient of the model’s output with respect to each input feature. The gradients are visualized to represent how variations in each input element, $X_{ij}$ (where $i,j$ are the spatial and temporal indices of $X$ , an input sample in $\mathbf{R}^{C\times T}$), influence the output prediction. The magnitude of each element $S_{ij}$ in $S$ illustrates the importance of the corresponding input pixel $X_{ij}$ to the output prediction.
To derive a spatial saliency map from these calculations, we first averaged $S$ across the temporal dimension to obtain $\bar{S}$. We then projected $\bar{S}$ onto the corresponding scalp electrode positions, creating a topographical map that illustrates the focal areas of brain activity relevant to the model’s decisions.
The saliency maps from the test set are plotted in Fig. 4, showing the averaged channel-wise importance across all subjects during the walking task. The results indicate that the highlighted EEG channels were predominantly concentrated in the central sensorimotor areas. Notably, electrodes such as Cz, C1, C2, CPz, CP1, and FC1 exhibited the highest saliency, suggesting their critical role in gait-related neural processing.
These observations reveal that the decoding model places strong emphasis on central regions typically associated with lower-limb motor control.
Figure 4: Saliency maps generated by NeuroDyGait computed across all subjects from the test set.
5.7 Ablation Studies
To understand the contribution of individual components in our framework, we conducted ablation experiments on GED and FBM dataset by selectively removing or modifying key loss terms and architectural modules. The results are summarized in Table 6.
These results collectively highlight the importance of each design component in our framework. In particular, the combination of contrastive pretraining with reconstruction and prediction. and dynamic domain fusion proves essential for learning transferable, motor-relevant EEG features and achieving strong decoding performance.
Table 6: Ablation study results on GED dataset.
Configuration Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
Full NeuroDyGait 0.6725 0.4458 0.7799
w/o Prediction Loss 0.6679 0.4345 0.7882
w/o Rel. Contrastive Loss 0.6617 0.4259 0.7942
w/o Reconstruction Loss 0.6551 0.4121 0.8038
w/o Stage I 0.6473 0.4164 0.8005
w/o Multi-head Fusion 0.6650 0.4280 0.7927
Table 7: Ablation study results on FBM dataset.
Configuration Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
Full NeuroDyGait 0.2585 0.0115 0.9952
w/o Prediction Loss 0.2515 0.0021 0.9999
w/o Rel. Contrastive Loss 0.2026 -0.1264 1.0612
w/o Reconstruction Loss 0.2444 -0.0075 1.0047
w/o Stage I 0.1924 -0.1180 1.0579
w/o Multi-head Fusion 0.2481 -0.0353 1.0182
6 Discussion
Our proposed framework, NeuroDyGait, demonstrates robust performance in decoding lower-limb kinematics from EEG across different subjects and datasets. By leveraging domain-invariant neural patterns, NeuroDyGait adapts well to varied populations and recording conditions, a critical advantage for clinical deployment.
Our findings further show that the pretrained EEG encoder learns transferable representations that can generalize across datasets. When pretrained on GED and transferred to FBM, the encoder provided a better initialization than random weights, improving convergence and predictive performance. This cross-domain robustness indicates that NeuroDyGait captures motor-relevant features that remain consistent across tasks and subjects, reducing the need for large dataset collection in every new setting.
Moreover, the fine-tuning results demonstrate NeuroDyGait’s capability to rapidly adapt to session-specific neural dynamics using minimal calibration data (less than 8% of a session). This efficient adaptation is particularly valuable for datasets with high inter-subject variability, such as FBM. It shows that NeuroDyGait not only captures generalizable cross-subject patterns but also effectively incorporates session-specific neural variations—an essential feature for handling nonstationarity in practical EEG applications.
Another key finding is the emergence of phase-specific structure in the EEG embeddings. t-SNE visualizations revealed that EEG representations naturally cluster according to gait phase, despite not being explicitly supervised for this attribute. This suggests that the contrastive learning objective in our phase-aware pretraining effectively captures the temporal dynamics inherent in gait cycles. Such phase-aware representations offer practical benefits in rehabilitation contexts, where identifying and targeting specific phases of the gait cycle can support real-time intervention and personalized therapy. For example, clinicians could leverage these representations to detect abnormalities or delays in particular gait phases and adjust therapy accordingly.
The domain relation-aware mechanism further strengthens the model’s robustness by adaptively focusing on the most informative source domains. The observed correlation between entropy and decoding accuracy suggests that confident and selective attention plays a critical role in enhancing prediction quality. Sessions with higher attention entropy—i.e., more dispersed domain attention—tend to have larger prediction errors. This mechanism is particularly beneficial in neurorehabilitation contexts, where subject-specific variability is high. The ability to dynamically adapt to varying source conditions ensures that the model maintains high performance across diverse individuals and sessions, aligning well with real-world clinical needs.
Spatial saliency analysis confirms that NeuroDyGait predominantly relies on motor-related cortical regions, particularly around the central midline such as Cz, C1, C2, CPz, CP1, and FC1. These electrodes correspond to the medial somatotopic representation of the lower limbs in the primary motor cortex and the primary somatosensory cortex, both of which are known to play crucial roles in gait execution  [seeber2015high, lotze1999activation, pfurtscheller1999event]. The alignment between these salient regions and established motor topography enhances the neurophysiological interpretability of NeuroDyGait, indicating that the model captures behaviorally meaningful neural signals rather than spurious noise. This targeted activation pattern not only strengthens confidence in the model’s decoding mechanism but also supports its applicability in motor-related BCI systems and neurorehabilitation. Channel-level saliency analysis provides a basis for designing more precise and personalized rehabilitation protocols. For example, monitoring saliency dynamics during recovery or different gait phases may allow clinicians to tailor stimulation or training strategies to specific motor areas  [hamedi2016electroencephalographic]. Overall, the central saliency profile suggests that NeuroDyGait leverages physiologically relevant information for both decoding and therapeutic applications.
Our ablation studies further corroborate the significance of key components in NeuroDyGait. Removing any one of the three losses used in Stage I resulted in decreased decoding performance, emphasizing the importance of each objective in shaping robust and informative EEG embeddings. Specifically, the reconstruction loss preserves the contextual information of motor dynamics, the prediction loss prevents the model from neglecting accurate prediction of the final motor state, and the relative contrastive loss enhances cross-modal alignment through stage-specific semantically structured embedding spaces. Furthermore, we removed the entire Stage I cross-modal pretraining and trained the model using only Stage II supervision. This led to a significant performance drop, which highlights that without the Biomechanical Supervision, the EEG encoder struggles to learn discriminative representations with motor semantics, underscoring the pivotal role of Stage I in bridging modalities and enhancing representational robustness. Additionally, the removal of multi-head fusion led to reduced performance, demonstrating the importance of selectively combining domain-specific information for robust generalization. These findings underscore the importance of each architectural choice in the proposed framework, reinforcing the model’s capacity to effectively learn and transfer motor-relevant neural patterns.
Although the current evaluation is limited to healthy individuals, the ability of NeuroDyGait to generalize across domains and datasets suggests its strong potential for use in clinical populations, including individuals with neurological impairments. Future work should focus on testing the model in diverse clinical settings, such as with patients who have suffered strokes or spinal cord injuries. Its ability to transfer across domains without requiring recalibration makes it an attractive solution for widespread use in clinical environments, where individual-specific models may not be feasible. Additionally, integrating multimodal data, such as combining EEG with other physiological signals, could further enhance its ability to monitor and adapt to the complex dynamics of motor recovery in real-time.
7 Conclusion
In this work, we introduced NeuroDyGait, a dual-stage and domain-generalization framework for decoding lower-limb kinematics from EEG. By combining a learnable cross-modal distance metric with relative contrastive learning, and a domain attention module for fusion-based decoding, the model demonstrates robust generalization across subjects and datasets without requiring subject-specific calibration. Extensive evaluations show that NeuroDyGait effectively captures transferable and temporally structured motor representations, as evidenced by its performance and interpretability in both time and spatial domains.
The integration of saliency analysis and contrastive embeddings provides further insight into the model’s internal representations, revealing alignment with neurophysiological principles of motor control. The identification of central motor regions as key contributors to decoding reinforces the relevance of the learned features, offering a basis for potential clinical translation.
While the current validation is limited to healthy individuals, the demonstrated cross-domain robustness suggests strong potential for extension to neurorehabilitation settings. Future work should focus on evaluating the model in clinical populations, exploring multimodal extensions, and examining real-time deployment scenarios. Overall, NeuroDyGait contributes to the growing body of research at the intersection of brain–computer interfaces and rehabilitative neuroscience by offering a scalable and interpretable approach for neural decoding in ambulatory motor tasks.
[…]